# Neil A. Chriss

### v1

@@ 5.2 Two firms trading in the presence of a non-strategic firm
Table 5 displays the type information for our first scenario. This example is as simple as can be. Each firm has two types and the types are virtually identical. All firms believe the market impact parameter $\kappa$ is equal to 1.0 and all firms trade to a target size of 1.0. The only difference among types is that firm 2’s type 1 believes there is a non-strategic trader trading to a target size of 5.0. We show below that even this difference profoundly impacts the strategies each type trades.
@@ Appendix B Acknowledgments
I extend my sincere thanks to the Machine Learning Research Group at Morgan Stanley who both originally suggested exploring Bayesian games and hosted my seminar on the topic during which they made quite a few important suggestions.

### v2

@@ 5.2 Two firms trading in the presence of a non-strategic firm
Table 6 displays the type information for our first scenario. This example is as simple as can be. Each firm has two types and the types are virtually identical. All firms believe the market impact parameter $\kappa$ is equal to 1.0 and all firms trade to a target size of 1.0. The only difference among types is that firm 2’s type 1 believes there is a non-strategic trader trading to a target size of 5.0 (as a reference this is the third element of $t_{2}^{1}$ in Table 6) below). We show below that even this difference profoundly impacts the strategies each type trades.
@@ Appendix B Acknowledgments
I extend my sincere thanks to the Machine Learning Research Group at Morgan Stanley who both originally suggested exploring Bayesian games and hosted my seminar on the topic during which they made quite a few important suggestions. For version two, a special thanks to Yuriy Nevmyvaka of Morgan Stanley’s for a careful reading of the first draft and identifying numerous typos which were corrected in the second draft.